# Nonequilibrium Probes of Quantum Geometry in Gapless Systems

### v1

@@ Appendix A Brief dictionary of CFT formulas
in terms of complexified generators $\ell_{m}$ satisfying the commutation relations $\ell_{m},\ell_{n}]=(m-n)\ell_{m+n}$ of the Witt algebra. [^12] This is the subalgebra $\operatorname{span}_{\mathbb{C}}(\{\ell_{n}\}_{n\in\mathbb{Z}})$ of the complexification $\operatorname{Vect}(S^{1})\otimes\mathbb{C}$. Note the isomorphism of the Witt algebra under $\ell_{n}\mapsto-\ell_{-n}$. This does not carry over to the Virasoro algebra (2.9) since its central term changes sign. Thus, Eq. (2.8) says that $L_{m}\equiv\mathfrak{u}\bigl(-\ell_{m}\bigr)+\frac{c}{24}\delta_{m,0}$ represents the (complexified) vector field $\ell_{m}$. Now, owing to the projective composition law (2.3), commutators of the operators (2.6) extend the algebra $\operatorname{Vect}(S^{1})$ by a central term:
$\big[\mathfrak{u}(\mathbf{X}),\mathfrak{u}(\mathbf{Y})\big]=\mathfrak{u}\big(-[\mathbf{X},\mathbf{Y}]\big)+ic\hskip 0.8pt\mathsf{c}(\mathbf{X},\mathbf{Y})\qquad\text{with}\qquad\mathsf{c}(\mathbf{X},\mathbf{Y})\equiv\oint\frac{\textrm{d}\varphi}{24\pi}X^{\prime}(\varphi)Y^{\prime\prime}(\varphi).$ (A.2)
@@
This $\mathsf{c}(\cdot,\cdot)$ is called the Gelfand–Fuchs cocycle; it is the algebra counterpart of the Bott cocycle (2.4). In fact, all nontrivial central extensions of $\operatorname{Vect}(S^{1})$ by $\mathbb{R}$ are isomorphic and generated by $\mathsf{c}(\cdot,\cdot)$ (see e.g. [4]), so the Virasoro central extension is unique up to redefinitions of generators. For the stress-energy tensor ${\mathcal{T}}(\theta)=i\mathfrak{u}(\delta(\varphi-\theta)\partial_{\varphi})$, Eq. (A.2) yields
@@
$\bigl[{\mathcal{T}}(\varphi),{\mathcal{T}}(\theta)\bigr]=-2i\delta^{\prime}(\varphi-\theta){\mathcal{T}}(\theta)+i\delta(\varphi-\theta){\mathcal{T}}^{\prime}(\theta)+\frac{c}{24\pi}i\delta^{\prime\prime\prime}(\varphi-\theta),$ (A.3)
@@
which are nothing but the Virasoro commutation relations (2.9) upon using Eq. (2.7). Lastly, we note that the Hermiticity $L_{m}^{\dagger}=L_{-m}^{\vphantom{\dagger}}$ follows directly from the unitarity of ${\cal U}(\cdot)$, or equivalently from the anti-Hermiticity of $\mathfrak{u}(\cdot)$.
Appendix B Virasoro quantum geometry away from the identity
@@
In this appendix, we briefly describe the quantum metric and the Berry curvature of a homogeneous parameter space (2.20) at points away from the identity. We then apply this to the case (2.12) that is relevant for Virasoro quantum geometry.
@@
Let $G$ be a Lie group, $\cal U$} a unitary representation of $G$, and consider the parameter space (2.20) for ‘deformed’ Hamiltonians of the form ${\cal U}(f)H_{0}\,{\cal U}(f)^{-1}$. We saw in Sec. 2 that the ensuing quantum metric and Berry curvature at the identity are given by Eq. (2.21). When evaluated at the coset $fG_{0}$ in the space (2.20), the metric and the curvature are given by
${\cal G}$ $=-\langle h|\mathfrak{u}(f^{-1}\textrm{d}f)\odot\mathfrak{u}(f^{-1}\textrm{d}f)|h\rangle+\langle h|\mathfrak{u}(f^{-1}\textrm{d}f)|h\rangle\odot\langle h|\mathfrak{u}(f^{-1}\textrm{d}f)|h\rangle,$ (B.1a)
${\cal F}$ $=2i\langle h|\mathfrak{u}(f^{-1}\textrm{d}f)\wedge\mathfrak{u}(f^{-1}\textrm{d}f)|h\rangle,$ (B.1b)
where $\mathfrak{u}(f^{-1}\textrm{d}f)$ is the representation of the left Maurer-Cartan form $f^{-1}\textrm{d}f$ on $G$. The latter is the map $f^{-1}\textrm{d}f:T_{f}G\to\mathfrak{g}$ such that, for any path $\gamma(t)$ in $G$ for which $\gamma(0)=f$, one has
$f^{-1}\textrm{d}f(\partial_{t}\gamma|_{t=0})\equiv\partial_{t}\big(f^{-1}\cdot\gamma(t)\big)\big|_{t=0}$ (B.2)
@@
For $G=\operatorname{Diff}(S^{1})$, the metric and the curvature at the identity are given by Eqs. (2.22)–(2.25). Let us now extend these expressions to the whole parameter space (2.12). Owing to Eq. (B.1), this is essentially achieved by replacing $\mathbf{X},\mathbf{Y}$ in Eqs. (2.24)–(2.25) by the left Maurer-Cartan form $f^{-1}\textrm{d}f$, and including tensor products where and when they are needed. Let us first discuss the Maurer-Cartan form in its own right. As briefly explained below Eq. (B.1), $f^{-1}\textrm{d}f$ is a Lie-algebra-valued one-form; so, for $G=\operatorname{Diff}(S^{1})$, it takes values in the space $\mathfrak{g}=\operatorname{Vect}(S^{1})$ of vector fields on the circle. Any tangent vector at $f\in\operatorname{Diff}(S^{1})$ can be seen as the time derivative $\partial_{t}\gamma_{t}|_{t=0}$ of a path $\gamma_{t}$ in $\operatorname{Diff}(S^{1})$ such that $\gamma_{0}=f$. This allows us to write the left Maurer-Cartan form on $\operatorname{Diff}(S^{1})$ by adapting Eq. (B.2) to a group operation given by the composition of functions:
$f^{-1}\textrm{d}f\big(\partial_{t}\gamma_{t}|_{t=0}\big)\equiv\partial_{t}\bigl(f^{-1}\circ\gamma_{t}\bigr)\big|_{t=0}=\frac{\partial_{t}\gamma_{t}(\varphi)|_{t=0}}{f^{\prime}(\varphi)}\partial_{\varphi}.$ (B.3)
@@
$f^{-1}\textrm{d}f\equiv\frac{\delta f(\varphi)}{f^{\prime}(\varphi)}\partial_{\varphi},$ (B.4)
@@
The compact notation (B.4) can now be combined with Eqs. (2.24)–(2.25) to express the quantum metric and the Berry curvature at any point $f$ in the group manifold $\operatorname{Diff}(S^{1})$ [or more precisely at any coset $f\circ S^{1}$ in the parameter space (2.12)]. Namely,
${\cal G}$ $=\oint\frac{\textrm{d}\varphi\,\textrm{d}\theta}{16\pi^{2}}\frac{(h{-}\frac{c}{24})\left(\left.\frac{\delta f}{f^{\prime}}\right|_{\varphi}{-}\left.\frac{\delta f}{f^{\prime}}\right|_{\theta}\right)^{2}+\frac{c}{24}\left(\left(\left.\frac{\delta f}{f^{\prime}}\right|_{\varphi}{-}\left.\frac{\delta f}{f^{\prime}}\right|_{\theta}\right)\!\Big.^{\prime}\,\right)^{2}}{\sin^{2}[(\theta-\varphi)/2]},$ (B.5a)
${\cal F}$ $=-2\oint\frac{\textrm{d}\varphi}{2\pi}\Big[\Big(h-\frac{c}{24}\Big)\frac{\delta f}{f^{\prime}}\wedge\Big(\frac{\delta f}{f^{\prime}}\Big)^{\prime}-\frac{c}{24}\frac{\delta f}{f^{\prime}}\wedge\Big(\frac{\delta f}{f^{\prime}}\Big)^{\prime\prime\prime}\Big],$ (B.5b)
where we use the notation $\delta f^{2}\equiv\delta f\odot\delta f$ for symmetrized products of one-forms. These expressions are explicit in principle, but complicated to use in practice. Again, the main lesson is that both of these tensor fields are fully determined by their value at the identity. In particular, we saw in Sec. 4 that the return probability of a slowly driven quantum state involves the quantum metric evaluated at a point $f$ far away from the identity, even though the actual formula involves the much simpler metric in Eqs. (2.22a) or (2.25) at the identity.
@@ Appendix C Linear response to Virasoro drives
The goal here is to derive the linear-response result (3.8) and its gradient expansion (3.11). The first step is to write the Hamiltonian in Eq. (3.2) as $H(t)=H_{0}+V_{\textrm{S}}(t)+O(\epsilon^{2})$ with
$V_{\textrm{S}}(t)\equiv\epsilon\bigl[\mathfrak{u}(\mathbf{X}_{t}),H_{0}\bigr]=\epsilon\frac{2\pi}{L}\bigl[\mathfrak{u}(\mathbf{X}_{t}),L_{0}\bigr]=i\epsilon\frac{2\pi}{L}\sum_{n\neq 0}n\hat{X}_{n}(t)L_{-n},$ (C.1)
@@
Since we wish to compute $\langle\phi(t)|i\mathfrak{u}(\mathbf{Y})|\phi(t)\rangle$ for any vector field $\mathbf{Y}$, it suffices to study the single-mode expectation value
$\langle\phi(t)|L_{-m}|\phi(t)\rangle$ (C.2)
for any given integer $m$. [Indeed, $i\mathfrak{u}(\mathbf{Y})$ is a linear combination of $L_{m}$s in Eq. (2.8).] To this end, we work in the interaction picture and define
$V_{\textrm{I}}(t)\equiv e^{iH_{0}t}V_{\textrm{S}}(t)e^{-iH_{0}t}=i\epsilon\frac{2\pi}{L}\sum_{n\neq 0}n\hat{X}_{n}(t)L_{-n}^{\textrm{I}}(t)=i\epsilon\frac{2\pi}{L}\sum_{n\neq 0}n\hat{X}_{n}(t)L_{-n}e^{i2\pi nt/L},$ (C.3)
with $L_{m}^{\textrm{I}}(t)\equiv e^{iH_{0}t}L_{m}e^{-iH_{0}t}=L_{m}e^{-i2\pi mt/L}$ obtained by solving $\partial_{t}L_{m}^{\textrm{I}}(t)=-i\frac{2\pi}{L}mL_{m}^{\textrm{I}}(t)$ with $L_{m}^{\textrm{I}}(0)=L_{m}$. Thus,
$\langle\phi(t)|L_{-m}|\phi(t)\rangle=\langle h|U_{\textrm{I}}(t)^{-1}L_{-m}^{\textrm{I}}(t)U_{\textrm{I}}(t)|h\rangle$ (C.4)
where the evolution operator $U_{\textrm{I}}(t)$ is expressible using the Dyson formula
$U_{\textrm{I}}(t)=I-i\int_{0}^{t}\textrm{d}s\,V_{\textrm{I}}(s)+O(\epsilon^{2}).$ (C.5)
This gives the Dyson series
@@
$=\langle h|L_{-m}|h\rangle e^{i2\pi mt/L}+\epsilon\sum_{n\neq 0}\int_{0}^{t}\textrm{d}s\,\chi_{m,n}(t,s)\hat{X}_{n}(s)+O(\epsilon^{2}),$ (C.6)
where we identified the Berry curvature ${\cal F}_{m,n}$ of Eq. (2.23b) and the susceptibility $\chi_{m,n}(t,s)$ of Eq. (3.9). The result (3.8) for operators $i\mathfrak{u}(\mathbf{Y})$ with Fourier components $\hat{Y}_{m}^{\phantom{*}}=\hat{Y}_{-m}^{*}$ follows.
@@
It remains to derive the gradient expansion (3.11). To this end, we use repeated integration by parts starting from Eq. (C.6), which yields
$\langle\phi(t)|L_{-m}|\phi(t)\rangle$ $=\langle h|L_{-m}|h\rangle e^{i2\pi mt/L}+\epsilon\sum_{n\neq 0}{\cal F}_{m,n}\biggl[\hat{X}_{n}(s)e^{i2\pi(mt+ns)/L}\biggr]_{s=0}^{t}$ (C.7)
@@
$R_{m}^{(N)}(t)=-\epsilon\sum_{n\neq 0}{\cal F}_{m,n}\biggl(-\frac{L}{2\pi in}\biggr)^{N}\int_{0}^{t}\textrm{d}s\,[\partial_{s}^{N+1}\hat{X}_{n}(s)]e^{i2\pi(mt+ns)/L}.$ (C.8)
@@
$\langle\phi(t)|i\mathfrak{u}(\mathbf{Y})|\phi(t)\rangle=\langle h|i\mathfrak{u}(\mathbf{Y})|h\rangle e^{i2\pi mt/L}+\epsilon\sum_{m,n\neq 0}\hat{Y}_{m}{\cal F}_{m,n}\left[\hat{X}_{n}(s)e^{i2\pi(mt+ns)/L}\right]_{s=0}^{t}$
$\qquad+\epsilon\sum_{m,n\neq 0}\hat{Y}_{m}{\cal F}_{m,n}\sum_{k=1}^{N}\biggl(-\frac{L}{2\pi in}\biggr)^{k}\biggl[\bigl[\partial_{s}^{k}\hat{X}_{n}(s)\bigr]e^{i2\pi(mt+ns)/L}\biggr]_{s=0}^{t}+\sum_{m\neq 0}\hat{Y}_{m}R_{m}^{(N)}(t)+O(\epsilon^{2}).$ (C.9)
This implies Eq. (3.11) since ${\cal F}_{m,n}$ is zero unless $m+n=0$.
Appendix D Rotating $\mathrm{SL}(2,\mathbb{R})$ drives
This appendix completes Sec. 5, for which it provides the explicit matrix expressions mentioned there.

### v2

@@ Appendix A Brief dictionary of CFT formulas
in terms of complexified generators $\ell_{m}$ satisfying the commutation relations $\ell_{m},\ell_{n}]=(m-n)\ell_{m+n}$ of the Witt algebra. [^18] This is the subalgebra $\operatorname{span}_{\mathbb{C}}(\{\ell_{n}\}_{n\in\mathbb{Z}})$ of the complexification $\operatorname{Vect}(S^{1})\otimes\mathbb{C}$. Note the isomorphism of the Witt algebra under $\ell_{n}\mapsto-\ell_{-n}$, which does not carry over to the Virasoro algebra (15) since the central term changes sign. The table’s nonchiral version of Eq. (14) thus says that $L^{\pm}_{m}\equiv\mathfrak{u}_{\pm}\bigl(-\ell_{\mp m}\bigr)+\frac{c}{24}\delta_{m,0}$ represents the (complexified) vector field $\ell_{m}$. Now, owing to the table’s nonchiral version of the projective composition law (9), commutators of the operators $\mathfrak{u}_{\pm}(\mathbf{X})$ extend the algebra $\operatorname{Vect}(S^{1})$ by a central term:
$\big[\mathfrak{u}_{\pm}(\mathbf{X}),\mathfrak{u}_{\pm}(\mathbf{Y})\big]=\mathfrak{u}_{\pm}\big({-}[\mathbf{X},\mathbf{Y}]\big)\pm ic\hskip 1.8pt\mathsf{c}(\mathbf{X},\mathbf{Y})$ (91)
@@
with $\mathsf{c}(\mathbf{X},\mathbf{Y})\equiv\oint\frac{\textrm{d}\varphi}{24\pi}X^{\prime}(\varphi)Y^{\prime\prime}(\varphi)$. The latter is the so-called Gelfand-Fuchs cocycle [81, 3]—the algebra counterpart of the Bott cocycle (10). In fact, all nontrivial central extensions of $\operatorname{Vect}(S^{1})$ by $\mathbb{R}$ are isomorphic and generated by $\mathsf{c}(\cdot,\cdot)$ (see e.g. [3]), so the Virasoro central extension is unique up to redefinitions of the generators. For the light-cone components ${\mathcal{T}}_{\pm}(\theta)=i\mathfrak{u}_{\pm}(\delta(\varphi-\theta)\partial_{\varphi})$ of the stress-energy tensor, Eq. (91) yields
@@
$\bigl[{\mathcal{T}}_{\pm}(\varphi),{\mathcal{T}}_{\pm}(\theta)\bigr]=\mp 2i\delta^{\prime}(\varphi-\theta)\,{\mathcal{T}}_{\pm}(\theta)\\ \pm i\delta(\varphi-\theta)\,{\mathcal{T}}_{\pm}^{\prime}(\theta)\pm\frac{c}{24\pi}\,i\,\delta^{\prime\prime\prime}(\varphi-\theta),$ (92)
@@
which are nothing but the Virasoro commutation relations (15) upon using the nonchiral version of Eq. (13) in the table. Lastly, note that Hermiticity $(L^{\pm}_{m})^{\dagger}=L^{\pm}_{-m}$ follows directly from the unitarity of ${\cal U}_{\pm}(\cdot)$, or equivalently from the anti-Hermiticity of $\mathfrak{u}_{\pm}(\cdot)$.
Appendix B Quantum geometry away from identity
@@
In this appendix, we briefly describe the quantum metric and the Berry curvature of a homogeneous parameter space (26) at points away from the identity. We then apply this to the case (18) that is relevant for Virasoro quantum geometry.
@@
Let $G$ be a Lie group, $\cal U$} a unitary representation of $G$, and consider the parameter space (26) for ‘deformed’ Hamiltonians of the form ${\cal U}(f)H_{0}\,{\cal U}(f)^{-1}$. We saw in Sec. II that the ensuing quantum metric and Berry curvature at the identity are given by Eq. (27). When evaluated at the coset $fG_{0}$ in the space (26), the metric and the curvature are given by
${\cal G}$ $=-\langle h|\mathfrak{u}(f^{-1}\textrm{d}f)\odot\mathfrak{u}(f^{-1}\textrm{d}f)|h\rangle$
$\quad\,+\langle h|\mathfrak{u}(f^{-1}\textrm{d}f)|h\rangle\odot\langle h|\mathfrak{u}(f^{-1}\textrm{d}f)|h\rangle,$ (93a)
${\cal F}$ $=2i\langle h|\mathfrak{u}(f^{-1}\textrm{d}f)\wedge\mathfrak{u}(f^{-1}\textrm{d}f)|h\rangle,$ (93b)
where $\odot$ and $\wedge$ are defined as in Eq. (25), and $\mathfrak{u}(f^{-1}\textrm{d}f)$ is the representation of the left Maurer-Cartan form $f^{-1}\textrm{d}f$ on $G$. The latter is the map $f^{-1}\textrm{d}f:T_{f}G\to\mathfrak{g}$ such that, for any path $\gamma(t)$ in $G$ for which $\gamma(0)=f$, one has
$f^{-1}\textrm{d}f(\partial_{t}\gamma|_{t=0})\equiv\partial_{t}\big(f^{-1}\cdot\gamma(t)\big)\big|_{t=0}$ (94)
@@
For $G=\operatorname{Diff}(S^{1})$, the metric and the curvature at the identity are given by Eqs. (29)–(32). Let us now extend these expressions to the whole parameter space (18). Owing to Eq. (93), this is essentially achieved by replacing $\mathbf{X},\mathbf{Y}$ in Eqs. (31)–(32) by the left Maurer-Cartan form $f^{-1}\textrm{d}f$, and including tensor products where and when they are needed. Let us first discuss the Maurer-Cartan form in its own right. As briefly explained below Eq. (93), $f^{-1}\textrm{d}f$ is a Lie-algebra-valued one-form; so, for $G=\operatorname{Diff}(S^{1})$, it takes values in the space $\mathfrak{g}=\operatorname{Vect}(S^{1})$ of vector fields on the circle. Any tangent vector at $f\in\operatorname{Diff}(S^{1})$ can be seen as the time derivative $\partial_{t}\gamma_{t}|_{t=0}$ of a path $\gamma_{t}$ in $\operatorname{Diff}(S^{1})$ such that $\gamma_{0}=f$. This allows us to write the left Maurer-Cartan form on $\operatorname{Diff}(S^{1})$ by adapting Eq. (94) to a group operation given by the composition of functions:
$f^{-1}\textrm{d}f\big(\partial_{t}\gamma_{t}|_{t=0}\big)\equiv\partial_{t}\bigl(f^{-1}\circ\gamma_{t}\bigr)\big|_{t=0}=\frac{\partial_{t}\gamma_{t}(\varphi)|_{t=0}}{f^{\prime}(\varphi)}\partial_{\varphi}.$ (95)
@@
$f^{-1}\textrm{d}f\equiv\frac{\delta f(\varphi)}{f^{\prime}(\varphi)}\partial_{\varphi},$ (96)
@@
The compact notation (96) can now be combined with Eqs. (31)–(32) to express the quantum metric and the Berry curvature at any point $f$ in the group manifold $\operatorname{Diff}(S^{1})$ [or more precisely at any coset $f\circ S^{1}$ in the parameter space (18)]. Namely,
${\cal G}$ $=\oint\frac{\textrm{d}\varphi\,\textrm{d}\theta}{16\pi^{2}}\frac{(h{-}\frac{c}{24})\Big(\left.\frac{\delta f}{f^{\prime}}\right|_{\varphi}{-}\left.\frac{\delta f}{f^{\prime}}\right|_{\theta}\Big)^{2}+\frac{c}{24}\Big(\Big(\left.\frac{\delta f}{f^{\prime}}\right|_{\varphi}{-}\left.\frac{\delta f}{f^{\prime}}\right|_{\theta}\Big)\!\Big.^{\prime}\,\Big)^{2}}{\sin^{2}[(\theta-\varphi)/2]},$ (97a)
${\cal F}$ $=-2\oint\frac{\textrm{d}\varphi}{2\pi}\Big[\Big(h-\frac{c}{24}\Big)\frac{\delta f}{f^{\prime}}\wedge\Big(\frac{\delta f}{f^{\prime}}\Big)^{\prime}-\frac{c}{24}\frac{\delta f}{f^{\prime}}\wedge\Big(\frac{\delta f}{f^{\prime}}\Big)^{\prime\prime\prime}\,\Big],$ (97b)
where we use the notation $(\delta f)^{2}\equiv\delta f\odot\delta f$ for symmetrized products of one-forms. These expressions are explicit in principle, but complicated to use in practice. Again, the main lesson is that both of these tensor fields are fully determined by their value at the identity. In particular, we saw in Sec. IV that the return probability of a slowly driven quantum state involves the quantum metric evaluated at a point $f$ far away from the identity, even though the actual formula involves the much simpler metric in Eqs. (29a) or (32) at the identity.
@@ Appendix C Linear response to Virasoro drives
Our goal here is to derive the linear-response result (39) and its gradient expansion (III.2). The first step is to write the Hamiltonian (33) as $H(t)=H_{0}+V_{\textrm{S}}(t)+O(\epsilon^{2})$ with
$V_{\textrm{S}}(t)\equiv\epsilon\bigl[\mathfrak{u}(\mathbf{X}_{t}),H_{0}\bigr]=i\epsilon\frac{2\pi}{L}\sum_{n\neq 0}n\hat{X}_{n}(t)L_{-n},$ (98)
@@
Since we wish to compute $\langle\phi(t)|i\mathfrak{u}(\mathbf{W})|\phi(t)\rangle$ for any vector field $\mathbf{W}$, it suffices to study the single-mode expectation value $\langle\phi(t)|L_{-m}|\phi(t)\rangle$ for any given integer $m$. [Indeed, $i\mathfrak{u}(\mathbf{W})$ is a linear combination of $L_{m}$s in Eq. (14).] To this end, we work in the interaction picture and define
$V_{\textrm{I}}(t)\equiv e^{iH_{0}t}V_{\textrm{S}}(t)e^{-iH_{0}t}\\ =i\epsilon\frac{2\pi}{L}\sum_{n\neq 0}n\hat{X}_{n}(t)L_{-n}^{\textrm{I}}(t)\\ =i\epsilon\frac{2\pi}{L}\sum_{n\neq 0}n\hat{X}_{n}(t)\,L_{-n}e^{i2\pi nt/L}$ (99)
with $L_{m}^{\textrm{I}}(t)\equiv e^{iH_{0}t}L_{m}e^{-iH_{0}t}=L_{m}e^{-i2\pi mt/L}$ obtained by solving $\partial_{t}L_{m}^{\textrm{I}}(t)=-i\frac{2\pi}{L}mL_{m}^{\textrm{I}}(t)$ with $L_{m}^{\textrm{I}}(0)=L_{m}$. Thus, $\langle\phi(t)|L_{-m}|\phi(t)\rangle=\langle h|U_{\textrm{I}}(t)^{-1}L_{-m}^{\textrm{I}}(t)U_{\textrm{I}}(t)|h\rangle$, where the evolution operator $U_{\textrm{I}}(t)$ is expressible using the Dyson formula $U_{\textrm{I}}(t)=I-i\int_{0}^{t}\textrm{d}s\,V_{\textrm{I}}(s)+O(\epsilon^{2})$. This gives the Dyson series
@@
$=\langle h|L_{-m}|h\rangle e^{i2\pi mt/L}+\epsilon\sum_{n\neq 0}\int_{0}^{t}\textrm{d}s\,\chi_{m,n}(t,s)\hat{X}_{n}(s)+O(\epsilon^{2}),$ (100)
where we identified the Berry curvature ${\cal F}_{m,n}$ of Eq. (30b) and the susceptibility $\chi_{m,n}(t,s)$ of Eq. (40). For operators $i\mathfrak{u}(\mathbf{W})$ with Fourier components $\hat{W}_{m}^{\phantom{*}}=\hat{W}_{-m}^{*}$, the result (39) follows.
@@
It remains to derive the gradient expansion (III.2). To this end, we use repeated integration by parts starting from Eq. (100), which yields
$\langle\phi(t)|L_{-m}|\phi(t)\rangle$ $=\langle h|L_{-m}|h\rangle e^{i2\pi mt/L}+\epsilon\sum_{n\neq 0}{\cal F}_{m,n}\biggl[\hat{X}_{n}(s)e^{i2\pi(mt+ns)/L}\biggr]_{s=0}^{t}$ (101)
@@
$R_{m}^{(N)}(t)=-\epsilon\sum_{n\neq 0}{\cal F}_{m,n}\biggl(-\frac{L}{2\pi in}\biggr)^{N}\int_{0}^{t}\textrm{d}s\,[\partial_{s}^{N+1}\hat{X}_{n}(s)]e^{i2\pi(mt+ns)/L}.$ (102)
@@
$\langle\phi(t)|i\mathfrak{u}(\mathbf{W})|\phi(t)\rangle=\langle h|i\mathfrak{u}(\mathbf{W})|h\rangle e^{i2\pi mt/L}+\epsilon\sum_{m,n\neq 0}\hat{W}_{m}{\cal F}_{m,n}\left[\hat{X}_{n}(s)e^{i2\pi(mt+ns)/L}\right]_{s=0}^{t}$
$\qquad+\epsilon\sum_{m,n\neq 0}\hat{W}_{m}{\cal F}_{m,n}\sum_{k=1}^{N}\biggl(-\frac{L}{2\pi in}\biggr)^{k}\biggl[\bigl[\partial_{s}^{k}\hat{X}_{n}(s)\bigr]e^{i2\pi(mt+ns)/L}\biggr]_{s=0}^{t}+\sum_{m\neq 0}\hat{W}_{m}R_{m}^{(N)}(t)+O(\epsilon^{2}).$ (103)
This implies Eq. (III.2) since ${\cal F}_{m,n}$ is zero unless $m+n=0$.
Appendix D Rotating $\mathrm{\bf SL}\bm{(}\bm{2}\bm{,}\mathbb{R}\bm{)}$ drives
This appendix completes Sec. V, for which it provides the explicit matrix expressions mentioned there.